1. Calculation of Cross-Rates. The spot rate of JPY with respect to GBP is 211.393 JPY/GBP.

The spot rate of GBP/EUR is 0.7909. a. Compute the JPY/EUR cross-rate. b. Compute the EUR/JPY cross-rate.
Answer #1

I GBP = 211-393 JPY 1 EUR = 0.7909 GBP 1 EUR = ? JPY 1 EUR = JPYX GBP D GBP EUR - 211-393 x 0.7909 11 EUR = 167. 1907 JPY)(6 EUR/JPY = EUR GB = 1 x 1 0.7909 211.393 1 Jpy = 0.0059819 EUR 1